A point charge q=2.00uC is located at the origin. Find the electric potential due to this charge at point x=4.0

1
Expert's answer
2022-03-28T14:24:53-0400

Electric potential

"V=\frac{kq}{r}"

"V=\frac{9\times10^9\times2\times10^{-6}}{4}=4.5\times10^{3}V"